40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 7 10:26:36 2018, keerthana, Update, PSL, observing the resonance signal corresponding to the injected frequency. photodetector_alignment.jpgdesign1.PNG
    Reply  Thu Jun 7 22:36:09 2018, not keerthana, Update, PSL, observing the resonance signal corresponding to the injected frequency. 
    Reply  Fri Jun 8 00:36:54 2018, gautam, Update, PSL, observing the resonance signal corresponding to the injected frequency. 
    Reply  Fri Jun 8 01:08:22 2018, johannes, Update, PSL, First light of AUX at YEND AG4395A_07-06-2018_205019.pdfPSL_AUX_SETUP.pdfAS_AUX_SETUP.pdf
       Reply  Mon Jun 18 14:32:42 2018, gautam, Update, PSL, Optics on AS table 
          Reply  Mon Jun 18 15:59:17 2018, johannes, Update, PSL, Optics on AS table 
             Reply  Mon Jun 18 16:57:54 2018, Koji, Update, PSL, Optics on AS table 
             Reply  Fri Jun 22 14:18:04 2018, Steve, Update, PSL, Optics on AS table 
Message ID: 13983     Entry time: Mon Jun 18 16:57:54 2018     In reply to: 13982
Author: Koji 
Type: Update 
Category: PSL 
Subject: Optics on AS table 

Of course, many (but no all) of the optics were custom-ordered back in ~2000.

ELOG V3.1.3-